OpenStudy (anonymous):

how do i write a in slope intercept form parallel to y= -3x+7 passing thru (-2,-5)

OpenStudy (anonymous):

Get the slope from the equation in the problem statement. It is in slope-intercept form which is y = mx + b. The form matches up and the slope is the value for "m". Take that "m" and use the point-slope formula:\[y - y _{1} = m(x - x _{1})\]where (x1, y1) is your point. Rearrange into the slope-intercept form which is already described earlier in this same post. Parallelism is achieved by using the same slope.